OpenStudy (anonymous):

What is the value of x if a rectangle has a perimeter of 50, a length of x+5, and width of 11.

OpenStudy (anonymous):

Will medal

OpenStudy (anonymous):

X is equal to 9

OpenStudy (anonymous):

Thanks, what equation did you use to get that

OpenStudy (anonymous):

2(x+5)+11=50 2x+16=50 2x+32=50 2x=18 x=9

OpenStudy (anonymous):

I was about to do it but Jeff done it for me Thanks Jeff

OpenStudy (anonymous):

Thanks to both of you